# A Survey of Lottery Ticket Hypothesis

### 07 Mar 2024 v1

@@ 6.1 Early Bird Tickets
The concept of Early-bird (EB) tickets is first introduced by(You et al., 2020). It refers to winning tickets identified in the early stages of training using cost-effective approaches like early stopping and low-precision training at high learning rates. This discovery aligns with prior observations that critical network patterns emerge early. Additionally, they introduce a mask distance metric for efficiently identifying EB tickets with minimal computational overhead, without requiring knowledge of the final winning tickets. By capitalizing on the existence of EB tickets and employing the proposed mask distance, we develop efficient training methods that involve identifying EB tickets through low-cost techniques and then continuing training exclusively on these EB tickets to reach the target accuracy. (Chen et al., 2020c)introduce EarlyBERT, drawing inspiration from the concept of Early-Bird Lottery Tickets proposed by (You et al., 2022b). The study put forth a computationally efficient training algorithm applicable to both pre-training and fine-tuning large-scale language models. By slimming down specific components of a transformer, the study identifies structured winning tickets in the early stages of BERT training, leading to efficient BERT training.
@@
Instead of pruning the trained networks, (Wang et al., 2020a) endeavors to prune networks during the initialization phase to save computational costs during training. They introduce a novel criterion known as Gradient Signal Preservation (GraSP). This method involves the removal of weights that have the least impact on reducing the gradient norm after pruning. (Rachwan et al., 2022) further refine the Gradient Signal Preservation (GraSP) by introducing Early Compression via Gradient Flow Preservation (EarlyCroP). This approach prioritizes the preservation of Gradient Flow (GF) and leverages the close relationship between the Neural Tangent Kernel (NTK) and GF. Consequently, it allows for network pruning with minimal disruption to the NTK, ensuring high levels of sparsity while maintaining performance comparable to that of the dense network. EarlyCroP requires only a single model training session and can be implemented either before or early in the training process.
@@ 6.3 Re-use winning tickets
Since the “retrainability” of a winning ticket is the most distinctive property of LTH, an intuitive method to reduce the cost is re-using the winning tickets across various tasks and models. (Morcos et al., 2019a) explore the potential of reusing winning tickets across different datasets and optimizers. They observe that winning tickets generated from larger datasets tend to transfer better. Moreover, these initializations also show promise across different optimization methods. These results indicate that winning ticket initializations from larger datasets possess inductive biases that benefit neural networks more broadly, offering promise for improved initialization methods. (Van Soelen & Sheppard, 2019) lay the foundation for a transfer learning technique that distills the original network to its crucial connections, obviating the need to freeze entire layers. the authors show that in the case of CNNs, networks can not only be pruned down to a mere 10% of their original parameters but these sparsified networks can also be re-trained on similar datasets with only a minor decrease in accuracy. Rooted in the LTH, this approach presents a promising alternative to conventional transfer learning methods and has shown encouraging initial outcomes (Burkholz et al., 2022) also theoretically prove the existence of universal lottery tickets. These tickets can be directly reused across various tasks without the need for retraining. (Desai et al., 2019) investigate the feasibility of re-training a sparse subnetwork obtained in one domain in a dissimilar domain and explore the impact of different initialization strategies during transfer. The experiments reveal that subnetworks obtained through lottery ticket training possess an inductive bias that transcends specific domains, enabling effective application across diverse domains.
@@
However, not all subnetworks exhibit universal transferability. (Chen et al., 2020a) investigate subnetworks extracted in the context of masked language modeling tasks, which demonstrate broad transferability. In contrast, subnetworks identified in other NLP tasks may possess limited applicability or even lack transferability altogether when applied in a transfer learning context. In addition to specific studies on the knowledge transfer of LTH in particular tasks and domains, there are ongoing investigations into the model’s capacity for continuous learning, known as lifelong learning. (Chen et al., 2020b) extend the LTH into the realm of lifelong learning, introducing the concept of "lifelong tickets." The research showcases that lifelong tickets can enhance learning performance across continual tasks. However, pruning networks in the lifelong setting presents significant challenges. Firstly, conducting greedy weight pruning in sequential tasks is hindered by a bias towards earlier tasks. Secondly, the compact network capacity of tickets in lifelong learning exacerbates the issue of catastrophic forgetting. To address these, two pruning options are proposed: top-down and bottom-up. The top-down approach extends iterative pruning over sequential tasks. In contrast, the bottom-up approach, allowing dynamic adjustment of model capacity, proves effective in mitigating early-stage excessive pruning. Additionally, the study introduces "lottery teaching," a method leveraging knowledge distillation with external unlabeled data to further counteract forgetting.
@@
(Burkholz et al., 2022) further substantiates that winning tickets can be applied across various tasks, suggesting a level of universality. This notion is formally established and theoretically validated, demonstrating the existence of universal tickets that do not necessitate additional training. Leveraging these validation results, they introduce several novel pruning approaches for strong lottery tickets and explicitly outline the sparse structure of the universal function family. Subsequently, (Huang et al., 2022) apply the LTH to CNNs integrating plug-and-play self-attention modules (SAMs). Their investigation reveals that complete integration of SAMs with all blocks may not lead to the maximal performance enhancement. Consequently, they introduce the concept of Lottery Ticket Hypothesis for Self-attention Networks, suggesting that a self-attention network encompasses a sparse subnetwork capable of expediting inference, reducing parameter overhead, and maintaining accuracy. This conjecture is substantiated by a combination of empirical observations and theoretical underpinnings.
@@ 6.4 Dual Lottery Ticket Hypothesis
Furthermore, LTH’s sparsification relies on the mask obtained from the pre-trained network, resulting in a fixed sparse structure. In contrast, DLTH depends on the randomly selected network, offering greater flexibility in the selection of sparse subnetworks. When compared to Pruning at Initialization (PI), DLTH focuses on transforming randomly selected subnetworks into trainable states, rather than starting from scratch to select a specific subnetwork. Thus, the sparsity structure in DLTH is controllable. In terms of accuracy, ResNet56 pruned using the RST method outperforms ResNet56 pruned using LTH at sparsity ratios of 50%, 70%, 90%, 95%, and 98% on the Cifar-10 and Cifar-100 dataset.
In summary, DLTH introduces a novel approach for training sparse networks that surpass existing methods (LTH and PI) in terms of generality, efficiency, accuracy, and flexibility in sparse structure selection. Inspired by the Lottery Ticket Hypothesis and Dual Lottery Ticket Hypothesis DLTH(Bai et al., 2022) (Wang et al., 2022a) propose two novel algorithms, UniLTH and UniDLTH, aimed at unifying LTH and DLTH.
UniLTH modifies the LTH approach by employing progressively increasing regularization on all weights of the neural network to discern their respective importance. On the other hand, UniDLTH applies to grow $L_{2}$ regularization exclusively to the masked weights for information extrusion. Both UniLTH and UniDLTH consist of two stages. In the first stage, the neural network is trained without any regularization, and early stopping is employed with a patience parameter, denoted as $\phi$. When the validation loss fails to decrease for $\phi$ consecutive epochs, training is halted, and parameters are reverted to their state from $x_{i}$ epochs earlier. In the second stage, iterative pruning is integrated with increased regularization to search or transform winning tickets. The network is alternatively trained with increased regularization and pruned until it attains the desired sparsity ratio.
@@ 7.3 Federated Learning
Federated learning (FL) is a collaborative decentralized learning paradigm to address privacy in machine learning. It suffers from heavy communication such as uploading and downloading large volumes of parameters in the model. LTH is introduced to reduce both the communication and computation costs through model compression (Itahara et al., 2020). The key idea is to obtain a winning ticket from the original model. LotteryFL (Li et al., 2020) further assign each client with a personalized winning ticket. Following LotteryFL, CELL (Seo et al., 2021) exploits downlink broadcast for communication efficiency and introduces a novel user grouping method to mitigate stragglers. FedLTN Mugunthan et al. (2022) proposes post-pruning without rewinding to achieve faster and greater model sparsity in FL. Babakniya et al. (2023) identify a challenge in federated learning (FL) on resource-limited edge nodes due to constrained computation and communication capabilities. They explore the use of off-the-shelf sparse learning algorithms, which train a binary sparse mask on each client, aiming to achieve a consistent sparse server mask with sparse weight tensors. However, their investigation reveals that this approach leads to a significant drop in accuracy compared to FL with dense models, particularly for clients with limited resources. They observe a notable lack of consensus among the trained sparsity masks on clients, hindering convergence for the server mask and potentially causing a substantial performance decline. In response, the authors propose "federated lottery aware sparsity hunting" (FLASH), a unified sparse learning framework designed to enable the server to attain a sparse sub-model, maintaining classification performance under highly resource-limited client settings. Additionally, they introduce "hetero-FLASH" to accommodate FL on different devices with varying parameter density requirements based on their resource constraints.
@@ 7.4 Reinforcement Learning
Due to the distributional shift inherent in reinforcement learning (RL) problems, the performance of winning lottery tickets can be affected (Vischer et al., 2022). Specifically, the feed-forward networks trained with behavioral cloning compared to reinforcement learning can be pruned to higher levels of sparsity, suggesting that the RL agents require more degrees of freedom. The analysis shows that the effect in learning paradigms can be attributed to the identified mask rather than the weight initialization. Also, the input layer mask selectively prunes entire input dimensions that turn out to be irrelevant to the task. The existence of winning tickets has also been investigated in a number of discrete-action space tasks, including both classic control and pixel control Yu et al. (2019). The result confirms that winning ticket initialization outperforms parameter-matched random initializations at extreme pruning rates in RL.
@@ 8.1 CV and NLP
Large-scale pre-trained language models such as transformers are usually over-parametrized Liang et al. (2021a). LTH has been applied to the GLUE benchmark, a collection of tools for evaluating the performance of models across a diverse set of existing natural language understanding tasks. These tasks include question-answering, sentiment analysis, and textual entailment, and an associated online platform for model evaluation, comparison, and analysis.
Recently, multimodal learning has been revolutionized by combining computer vision and natural language processing technologies. An empirical study has been performed to assess whether winning tickets also exist in pre-trained vision-and-language (VL) models Gan et al. (2022). The authors consolidate 7 representative VL tasks: visual question answering, visual commonsense reasoning, visual entailment, referring expression comprehension, image-text retrieval, GQA, and NLVR. It turns out that the sparse model can not strictly match the performance of the full model. However, a relaxed” winning tickets at 50%-70% sparsity maintains 99% of the full accuracy.
@@ 9.1 Acceleration in Practice
Although diffusion models have shown impressive performance in capturing distributions and sample quality, they are notoriously slow to generate data due to the long chain of reversing the diffusion process. It is intuitive to mitigate the computational cost by pruning the reverse model. An open question is can we use sub-networks with different sparsity in the reverse process? Since a winning ticket can be considered an equivalent version of the original model, we can leverage a sparser sub-network in the late stage of denoising to further improve efficiency. Intuitively, the noise in the later reverse process has been reduced and it will be easier to optimize. The challenge lies in how to optimize a dense model and a winning ticket while guaranteeing the convergence of training. Another challenge is to decide at which step to use the winning ticket. We leave this open question for future investigation which we hope to stimulate the research on improving the efficiency of the diffusion model.

### 12 Mar 2024 v2

@@ 6.1 Early Bird Tickets
The concept of Early-bird (EB) tickets is first introduced by You et al. (2020). It refers to winning tickets identified in the early stages of training using cost-effective approaches like early stopping and low-precision training at high learning rates. This discovery aligns with prior observations that critical network patterns emerge early. Additionally, they introduce a mask distance metric for efficiently identifying EB tickets with minimal computational overhead, without requiring knowledge of the final winning tickets. By capitalizing on the existence of EB tickets and employing the proposed mask distance, we develop efficient training methods that involve identifying EB tickets through low-cost techniques and then continuing training exclusively on these EB tickets to reach the target accuracy. Chen et al. (2020c) introduce EarlyBERT, drawing inspiration from the concept of Early-Bird Lottery Tickets proposed by (You et al., 2022b). The study put forth a computationally efficient training algorithm applicable to both pre-training and fine-tuning large-scale language models. By slimming down specific components of a transformer, the study identifies structured winning tickets in the early stages of BERT training, leading to efficient BERT training.
@@
Instead of pruning the trained networks, (Wang et al., 2020a) endeavor to prune networks during the initialization phase to save computational costs during training. They introduce a novel criterion known as Gradient Signal Preservation (GraSP). This method involves the removal of weights that have the least impact on reducing the gradient norm after pruning. (Rachwan et al., 2022) further refine the Gradient Signal Preservation (GraSP) by introducing Early Compression via Gradient Flow Preservation (EarlyCroP). This approach prioritizes the preservation of Gradient Flow (GF) and leverages the close relationship between the Neural Tangent Kernel (NTK) and GF. Consequently, it allows for network pruning with minimal disruption to the NTK, ensuring high levels of sparsity while maintaining performance comparable to that of the dense network. EarlyCroP requires only a single model training session and can be implemented either before or early in the training process.
@@ 6.3 Re-use winning tickets
Since the “retrainability” of a winning ticket is the most distinctive property of LTH, an intuitive method to reduce the cost is re-using the winning tickets across various tasks and models. (Morcos et al., 2019a) explore the potential of reusing winning tickets across different datasets and optimizers. They observe that winning tickets generated from larger datasets tend to transfer better. Moreover, these initializations also show promise across different optimization methods. These results indicate that winning ticket initializations from larger datasets possess inductive biases that benefit neural networks more broadly, offering promise for improved initialization methods. (Van Soelen & Sheppard, 2019) lay the foundation for a transfer learning technique that distills the original network to its crucial connections, obviating the need to freeze entire layers. The authors show that in the case of CNNs, networks can not only be pruned down to a mere 10% of their original parameters but these sparsified networks can also be re-trained on similar datasets with only a minor decrease in accuracy. Rooted in the LTH, this approach presents a promising alternative to conventional transfer learning methods and has shown encouraging initial outcomes. Burkholz et al. (2022) also theoretically prove the existence of universal lottery tickets. These tickets can be directly reused across various tasks without the need for retraining. (Desai et al., 2019) investigate the feasibility of re-training a sparse subnetwork obtained from one domain to a dissimilar domain and explores the impact of different initialization strategies during transfer. The experiments reveal that subnetworks obtained through lottery ticket training possess an inductive bias that transcends specific domains, enabling effective application across diverse domains.
@@
However, not all subnetworks exhibit universal transferability. (Chen et al., 2020a) investigate subnetworks extracted in the context of masked language modeling tasks, which demonstrate broad transferability. In contrast, subnetworks identified in other NLP tasks may possess limited applicability or even lack transferability altogether when applied in a transfer learning context. In addition to specific studies on the knowledge transfer of LTH in particular tasks and domains, there are ongoing investigations into the model’s capacity for continuous learning, known as lifelong learning. (Chen et al., 2020b) extend the LTH into the realm of lifelong learning, introducing the concept of "lifelong tickets." This research showcases that lifelong tickets can enhance learning performance across continual tasks. However, pruning networks in the lifelong setting presents significant challenges. Firstly, conducting greedy weight pruning in sequential tasks is hindered by a bias towards earlier tasks. Secondly, the compact network capacity of tickets in lifelong learning exacerbates the issue of catastrophic forgetting. To address these, two pruning options are proposed: top-down and bottom-up. The top-down approach extends iterative pruning over sequential tasks. In contrast, the bottom-up approach, allowing dynamic adjustment of model capacity, proves effective in mitigating early-stage excessive pruning. Additionally, the study introduces "lottery teaching," a method leveraging knowledge distillation with external unlabeled data to further counteract forgetting.
@@
(Burkholz et al., 2022) further substantiates that winning tickets can be applied across various tasks, suggesting a level of universality. This notion is formally established and theoretically validated, demonstrating the existence of universal tickets that do not necessitate additional training. Leveraging these validation results, they introduce several novel pruning approaches for strong lottery tickets and explicitly outline the sparse structure of the universal function family. Subsequently, (Huang et al., 2022) apply the LTH to CNNs integrating plug-and-play self-attention modules (SAMs). Their investigation reveals that complete integration of SAMs with all blocks may not lead to maximal performance enhancement. Consequently, they introduce the concept of Lottery Ticket Hypothesis for Self-attention Networks, suggesting that a self-attention network encompasses a sparse subnetwork capable of expediting inference, reducing parameter overhead, and maintaining accuracy. This conjecture is substantiated by a combination of empirical observations and theoretical underpinnings.
@@ 6.4 Dual Lottery Ticket Hypothesis
Furthermore, LTH’s sparsification relies on the mask obtained from the pre-trained network, resulting in a fixed sparse structure. In contrast, DLTH depends on the randomly selected network, offering greater flexibility in the selection of sparse subnetworks. When compared to Pruning at Initialization (PI), DLTH focuses on transforming randomly selected subnetworks into trainable states, rather than starting from scratch to select a specific subnetwork. Thus, the sparsity structure in DLTH is controllable. In terms of accuracy, ResNet56 pruned using the RST method outperforms ResNet56 pruned using LTH at sparsity ratios of 50%, 70%, 90%, 95%, and 98% on the Cifar-10 and Cifar-100 datasets.
In summary, DLTH introduces a novel approach for training sparse networks that surpass existing methods (LTH and PI) in terms of generality, efficiency, accuracy, and flexibility in sparse structure selection. Inspired by the Lottery Ticket Hypothesis and Dual Lottery Ticket Hypothesis DLTH (Bai et al., 2022), (Wang et al., 2022a) propose two novel algorithms, UniLTH and UniDLTH, aimed at unifying LTH and DLTH. UniLTH modifies the LTH approach by employing progressively increasing regularization on all weights of the neural network to discern their respective importance. On the other hand, UniDLTH applies to grow $L_{2}$ regularization exclusively to the masked weights for information extrusion. Both UniLTH and UniDLTH consist of two stages. In the first stage, the neural network is trained without any regularization, and early stopping is employed with a patience parameter, denoted as $\phi$. When the validation loss fails to decrease for $\phi$ consecutive epochs, training is halted, and parameters are reverted to their state from $x_{i}$ epochs earlier. In the second stage, iterative pruning is integrated with increased regularization to search or transform winning tickets. The network is alternatively trained with increased regularization and pruned until it attains the desired sparsity ratio.
@@ 7.3 Federated Learning
Federated learning (FL) is a collaborative decentralized learning paradigm to address privacy in machine learning. It suffers from heavy communication such as uploading and downloading large volumes of parameters in the model. LTH is introduced to reduce both the communication and computation costs through model compression (Itahara et al., 2020). The key idea is to obtain a winning ticket from the original model. LotteryFL (Li et al., 2020) further assign each client with a personalized winning ticket. Following LotteryFL, CELL (Seo et al., 2021) exploits downlink broadcast for communication efficiency and introduces a novel user grouping method to mitigate stragglers. FedLTN (Mugunthan et al., 2022) proposes post-pruning without rewinding to achieve faster and greater model sparsity in FL. Babakniya et al. (2023) identify a challenge in federated learning (FL) on resource-limited edge nodes due to constrained computation and communication capabilities. They explore the use of off-the-shelf sparse learning algorithms, which train a binary sparse mask on each client, aiming to achieve a consistent sparse server mask with sparse weight tensors. However, their investigation reveals that this approach leads to a significant drop in accuracy compared to FL with dense models, particularly for clients with limited resources. They observe a notable lack of consensus among the trained sparsity masks on clients, hindering convergence for the server mask and potentially causing a substantial performance decline. In response, the authors propose "federated lottery aware sparsity hunting" (FLASH), a unified sparse learning framework designed to enable the server to attain a sparse sub-model, maintaining classification performance under highly resource-limited client settings. Additionally, they introduce "hetero-FLASH" to accommodate FL on different devices with varying parameter density requirements based on their resource constraints.
@@ 7.4 Reinforcement Learning
Due to the distributional shift inherent in reinforcement learning (RL) problems, the performance of winning lottery tickets can be affected (Vischer et al., 2022). Specifically, the feed-forward networks trained with behavioral cloning compared to reinforcement learning can be pruned to higher levels of sparsity, suggesting that the RL agents require more degrees of freedom. The analysis shows that the effect in learning paradigms can be attributed to the identified mask rather than the weight initialization. Also, the input layer mask selectively prunes entire input dimensions that turn out to be irrelevant to the task. The existence of winning tickets has also been investigated in a number of discrete-action space tasks, including both classic control and pixel control (Yu et al., 2019). The result confirms that winning ticket initialization outperforms parameter-matched random initializations at extreme pruning rates in RL.
@@ 8.1 CV and NLP
Large-scale pre-trained language models such as transformers are usually over-parametrized (Liang et al., 2021a). LTH has been applied to the GLUE benchmark, a collection of tools for evaluating the performance of models across a diverse set of existing natural language understanding tasks. These tasks include question-answering, sentiment analysis, and textual entailment, and an associated online platform for model evaluation, comparison, and analysis.
Recently, multimodal learning has been revolutionized by combining computer vision and natural language processing technologies. An empirical study has been performed to assess whether winning tickets also exist in pre-trained vision-and-language (VL) models Gan et al. (2022). The authors consolidate 7 representative VL tasks: visual question answering, visual commonsense reasoning, visual entailment, referring expression comprehension, image-text retrieval, GQA, and NLVR. It turns out that the sparse model can not strictly match the performance of the full model. However, a “relaxed” winning ticket at 50%-70% sparsity can reach 99% of the full accuracy.
@@ 9.1 Acceleration in Practice
Although diffusion models have shown impressive performance in capturing distributions and sample quality, they are notoriously slow to generate data due to the long chain of reversing the diffusion process. It is intuitive to mitigate the computational cost by pruning the reverse model. An open question remains: Can we use sub-networks with different sparsity in the reverse process? Since a winning ticket can be considered an equivalent version of the original model, we can leverage a sparser sub-network in the late stage of denoising to further improve efficiency. Intuitively, the noise in the later reverse process has been reduced and it will be easier to optimize. The challenge lies in how to optimize a dense model and a winning ticket while guaranteeing the convergence of training. Another challenge is to decide at which step to use the winning ticket. We leave this open question for future investigation which we hope will stimulate the research on improving the efficiency of the diffusion model.